Alice decides to set up an RSA public key encryption using the two primes p= 31 and p= 41 and the encryption key e= 11.You must show all calculations, including MOD-calculations using the division algorithm!
Bob decides to send the message M= 30to her using this encryption. What is the code C that he will send her?
What is Alice’s decryption key d?
Alice also receives the message C= 101from Carla. What was her original message M?
1
Expert's answer
2020-12-13T18:21:35-0500

"n = p \times q \\

n = 31 \times 41 = 1271 \\

Euler's\; totient\; of \;n \\

ɸ(n) = (p-1)(q-1) \\

= 30 \times 40 = 1200 \\

gcd(ɸ(n),e) = gcd(1200,11) = 1 \\

1< e<ɸ(n) \\

edmodɸ(n) = 1 \\

= ed \equiv modɸ(n) \\

d = e^{-1}modɸ(n)

11dmod1200 = 1 \\

x mod y \equiv \frac{x}{y} \equiv R \\

\frac{11d}{1200} \\

R = 1 \\

d = 1091"

decryption key d = 1091

Encrypting message M with public key (n,e)

"C = M^emodn \\

= 30^{ʹʹ}mod1200 \\

C = 557"

For C = 101 Decrypting with private key (1271, 1091)

"M = C^dmod n \\

= 101^{1091}mod1271 = 95 \\

d = \frac{1 + iɸ(n)}{e}"

If i = 1

"d= \frac{1 + 1200}{11} = 109.18"

If i = 2

"d= \frac{1 + 2400}{11} = 218.27"

If i = 10

"d= \frac{1 + 12000}{11} = 1091"